OpenStudy (anonymous):

Xian saw a blouse that costs $37.89,including tax. The cost of materials to make the blouse is $20.98,plus tax. The tax on all the materials that she needs is 8.25%. About how much will she save if she makes the blouse?

OpenStudy (anonymous):

Whats the tax on the blouse to begin with?

OpenStudy (anonymous):

it is 8.25%

OpenStudy (anonymous):

You don't know. You only know the tax of the materials.

OpenStudy (anonymous):

Approx. $18.31

OpenStudy (anonymous):

my choices are $15.27 $15.81 $15.18

OpenStudy (anonymous):

Whats the sales tax were u live?

OpenStudy (anonymous):

Total value of blouse materials= (20.89 * 0.0825) + 20.89 =22.71 37.89 - 22.71 = 15.18

OpenStudy (anonymous):

thanks romero..!!:)) ur such a lifesavor!!!